Solve system by substitution or addition
2y-X=3
x=3y-5

Answer by oscargut(2103)

2y-X=3
x=3y-5
if we do the substitution of x in the firs eq
2y-(3y-5)=3
-y+5=3
-y=-2
y=2
and x=3(2)-5=1
Answer: x=1,y=2